The amount of CO2 in the ocean is approximately 280 ppm. What is the mole fraction of CO2 in a liter of ocean water?
1
Expert's answer
2020-08-29T14:01:15-0400

Dencity of ocean water is 1027 kg/m3

280 ppm of CO2 is 280 mg/l and approximately 280 mg/kg

M (CO2) = 44 g/mol

Moles/kg of CO2 = 280 mg/kg / 44 g/mol / 1000 = 6.36 x 10-3 mol/kg CO2

Molarity = 6.36 x 10-3 mol/kg CO2 x 1027 kg/l = 6.53 M

Mass of ocean water is 1000 х 1027 kg/m3 = 1 027 000 g

M (H2O) = 18 g/mol

Moles of ocean water are 1 027 000 g / 18 g/mol = 57 056 moles

Mole fraction of CO2 in ocean water is: 6.53 / (6.53+57 056) = 0.0001 = 1 x 10-4